OpenStudy (anonymous):

Can anyone help me with “Using the Rules of Exponents? If so, please show me the steps. I have two equations. Thank you. 27 ^-2/3÷27 ^2-1/3 Here is the second equation: (a ½b) ½ (ab ½)

OpenStudy (whpalmer4):

Rule for division: \[\frac{x^n}{x^m} = x^{n-m}\]

OpenStudy (anonymous):

Hi WH. Can you show me the steps?

OpenStudy (whpalmer4):

yeah. take the exponent from your numerator. subtract the exponent from the denominator. that's the exponent of the result. \[\frac{x^4}{x^1} = x^{4-1} = x^3\]

OpenStudy (whpalmer4):

it's pretty easy to see for integer exponents why this is true: \[\frac{x^4}{x^1} = \frac{x*x*x*x}{x} = \frac{\cancel{x}*x*x*x}{\cancel{x}} = x*x*x = x^3\]

OpenStudy (anonymous):

Is this for equation (a ½b) ½ (ab ½)? If so why are you using z?

OpenStudy (anonymous):

WH are you there?

OpenStudy (whpalmer4):

I'm showing the pattern of how to do these problems. you'll have to apply the pattern to your problem.

OpenStudy (jhannybean):

\[(a\times \frac12 b)\times \frac12 \times (ab\cdot \frac12)\]???

OpenStudy (jhannybean):

are the 1/2 s powers?...or...

OpenStudy (anonymous):

They are exponents.

OpenStudy (jhannybean):

like this? \[\large (a^{1/2}b)\times (ab^{1/2})\]?

OpenStudy (jhannybean):

It's a little confusing since you don't have a carrot "^" indicating whether its an exponent or a multiplier.

OpenStudy (anonymous):

Yes, and there is a ^1/2 in the center of the two (a ½b) ½ (ab ½).

OpenStudy (anonymous):

I wrote it in word first using the superscript! For some reason it did not show up in this format, I will have to remember that.

OpenStudy (jhannybean):

Oh okay. \[\large (a^{1/2}b)^{1/2}\times (ab^{1/2})\]

OpenStudy (anonymous):

Yes, can you show me the steps?

OpenStudy (jhannybean):

Yeah, it'd be wise. The person helping you will get confused otherwise, haha.

OpenStudy (jhannybean):

Take the first group, \(\large( a^{1/2}b)^{1/2}\) multiply in the 1/2 power to both of these variables, what do you get?

OpenStudy (anonymous):

a^1/2 b^1/2

OpenStudy (jhannybean):

Nope. \[\large (a^{1/2 \cdot 1/2})(b^{1\cdot 1/2}) =?\]

OpenStudy (anonymous):

1

OpenStudy (jhannybean):

No.

OpenStudy (anonymous):

ab

OpenStudy (jhannybean):

What do the exponents equal?

OpenStudy (anonymous):

TH

OpenStudy (anonymous):

The number or variable exponentially.

OpenStudy (jhannybean):

Both.... What is \(\large \frac{1}{2} \cdot \frac{1}{2} = ?\)

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

Two halve = 1

OpenStudy (whpalmer4):

not adding, multiplying! 1/2 of 1/2 is...

OpenStudy (anonymous):

WH if I do not know the answer to this questions which is why I asked for the steps. I am new to algebra.

OpenStudy (whpalmer4):

the center dot is a sign for multiplication. do you know how to multiply fractions?

OpenStudy (anonymous):

I have learned that multiplying negative exponents is different since we sometimes us the reciprocal which confuses me.

OpenStudy (anonymous):

I know that th

OpenStudy (anonymous):

There is no negative in my second question but there is in the first one so I asked for someone to walk me through the steps.

OpenStudy (whpalmer4):

that's fine. but what is 1/2 * 1/2?

OpenStudy (anonymous):

1

OpenStudy (whpalmer4):

no. 1/2 + 1/2 = 1. what is 1/2 of 1/2?

OpenStudy (anonymous):

2/2

OpenStudy (whpalmer4):

oh dear... if you have 1 of something, and you divide it to two equal pieces, each piece is how much?

OpenStudy (anonymous):

each is 1/2

OpenStudy (whpalmer4):

okay, and if you divide each of those pieces into two equal pieces, how much are they?

OpenStudy (anonymous):

1/3

OpenStudy (whpalmer4):

so 1 thing divided into 4 equal pieces makes the pieces be 1/3 each?

OpenStudy (jhannybean):

no that is taking 1 piece and dividing it into 3 equal parts.each piece would then be 1/3 .

OpenStudy (whpalmer4):

what is 1/2? as a decimal?

OpenStudy (anonymous):

.5

OpenStudy (whpalmer4):

okay. what is that, divided by 2?

OpenStudy (jhannybean):



OpenStudy (jhannybean):